OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1).

OpenStudy (tkhunny):

Rise = Change in y = -6 - 1 Run= Change in x = -6 - (-3) Where else will this go?

OpenStudy (anonymous):

it would become -7/-3?

OpenStudy (tkhunny):

Or just 7/3

OpenStudy (anonymous):

ok, thanks! :)